sasa:

) Find the exact value: cos(7pi/12) Part I: Find two common angles that either add up to or differ by 7pi/12. Rewrite this problem as the cosine of either a sum or a difference of those two angles. Part II: Evaluate the expression. 2) tan(17pi/12) Part I: Find two common angles that either add up to or differ by 17pi/12. Part II: Evaluate. 3) sin(195 degrees) Part I: pretty much the same as above Part II: Evaluate. 4) cot(15 degrees) Part I: Find two common angles that differ by 15 degrees. Rewrite this problem as the cotangent of a difference of those two angles. Part II: Evaluate the expression.

surjithayer:

do you know how to do it?

surjithayer:

1. \[\frac{ 7 \pi }{12 }=\frac{ \pi }{ 3 }+\frac{ \pi }{ 4 }\] \[\cos (\frac{ 7\pi }{ 12})=\cos (\frac{ \pi }{ 3 }+\frac{ \pi }{ 4})\] \[=\cos \frac{ \pi }{ 3 }\cos \frac{ \pi }{ 4 }-\sin \frac{ \pi }{ 3 }\sin \frac{ \pi }{ 4}\] =?

surjithayer:

2. \[\frac{ 17 \pi }{12 }=2\pi-\frac{ 7\pi }{ 12 }\]

surjithayer:

\[\tan (\frac{ 17\pi }{ 12 })=\tan (2\pi-\frac{ 7\pi }{ 12})=-\tan (\frac{ 7 \pi }{12})\]

surjithayer:

2. \[=-\tan (\frac{ \pi }{ 3 }+\frac{ \pi }{ 4})\] \[=-\frac{ \tan \frac{ \pi }{ 3 }+\tan \frac{ \pi }{ 4} }{ 1-\tan \frac{ \pi }{ 3 } \tan \frac{ \pi }{ 4 }}\] =? 3. \[\sin (195)=\sin (180+15)=-\sin 15=-\sin (45-30)\] \[=-(\sin45\cos30-\cos45\sin30)=?\]

surjithayer:

4. \[\cot15=\cot(90-75)=\tan 75\] \[=\tan (45+30)\] \[=\frac{ \tan 45+\tan30 }{ 1-\tan 45 \tan 30 }=?\] or 4. \[\cot15=\frac{ 1 }{ \tan 15 }=\frac{ 1 }{ \tan (45-30)}\] \[=\frac{ 1 }{\frac{ \tan45-\tan30 }{ 1+\tan 45\tan30 } }\] \[=\frac{ 1+\tan45 \tan 30 }{ \tan45-\tan30 }=?\]